### v1

# Med3DVLM: An Efficient Vision-Language Model for 3D Medical Image Analysis

Abstract
Vision-language models (VLMs) have shown promise in 2D medical image analysis, but extending them to 3D remains challenging due to the high computational demands of volumetric data and the difficulty of aligning 3D spatial features with clinical text. We present Med3DVLM, a 3D VLM designed to address these challenges through three key innovations: (1) DCFormer, an efficient encoder that uses decomposed 3D convolutions to capture fine-grained spatial features at scale; (2) SigLIP, a contrastive learning strategy with pairwise sigmoid loss that improves image-text alignment without relying on large negative batches; and (3) a dual-stream MLP-Mixer projector that fuses low- and high-level image features with text embeddings for richer multi-modal representations. We evaluate our model on the M3D dataset, which includes radiology reports and VQA data for 120,084 3D medical images. Results show that Med3DVLM achieves superior performance across multiple benchmarks. For image-text retrieval, it reaches 61.00% R@1 on 2,000 samples, significantly outperforming the current state-of-the-art M3D-LaMed model (19.10%). For report generation, it achieves a METEOR score of 36.42% (vs. 14.38%). In open-ended visual question answering (VQA), it scores 36.76% METEOR (vs. 33.58%), and in closed-ended VQA, it achieves 79.95% accuracy (vs. 75.78%). These results highlight Med3DVLM’s ability to bridge the gap between 3D imaging and language, enabling scalable, multi-task reasoning across clinical applications. Our code is publicly available at https://github.com/mirthAI/Med3DVLM.
11footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
1 Introduction
Medical image analysis plays a crucial role in diagnosing and treating diseases such as cancer, cardiovascular conditions, and neurological disorders. However, existing models are often task-specific, lack adaptability to new tasks, and do not support real-time user interactions. Vision-language models (VLMs), such as Contrastive Language-Image Pre-Training (CLIP)[27] and Large Language and Vision Assistant (LLaVA)[19], offer greater versatility in dynamic clinical settings by aligning medical images with textual reports. CLIP leverages contrastive learning on image-text pairs, enabling zero-shot image classification and image-text retrieval, while LLaVA extends CLIP by integrating its visual encoder with a large language model (LLM), facilitating interactive tasks such as report generation and visual question answering (VQA). Despite their success in analyzing 2D medical images, such as chest X-rays[8] and pathology slides[21], their application to 3D imaging remains limited.
3D imaging modalities, such as computed tomography (CT) and magnetic resonance imaging (MRI), provide volumetric data that capture spatial details unavailable in 2D images. However, extending VLMs to 3D presents significant challenges. First, a 3D scan consists of hundreds of slices, making slice-by-slice analysis prone to losing global context. Second, directly building 3D VLMs significantly increases computational complexity due to the higher dimensionality. Third, the scarcity of publicly available 3D image-report pairs further limits model development.
Several VLMs have attempted to bridge this gap, but each has limitations in the 3D setting. For example, PMC-CLIP[17], trained on large-scale biomedical literature images, is restricted to 2D inputs, leading to poor performance on 3D image understanding. RadFM[38] unifies 2D and 3D data but is primarily optimized for text generation tasks, such as VQA, and struggles with broader image-text understanding. More recently, M3D-LaMed[2] was introduced as a generalist multi-modal model for 3D medical image analysis. By combining a CLIP-pretrained 3D visual encoder with a 3D spatial pooling perceiver, M3D-LaMed enables direct reasoning on 3D scans and achieves state-of-the-art performance across multiple benchmarks, including image-text retrieval, report generation, open-ended and closed-ended VQA, as well as 3D segmentation and localization tasks.
Despite these advancements, M3D-LaMed has notable limitations. Its 3D backbone, similar to many other 3D VLMs, incurs high computational costs when processing high-resolution 3D volumes, as standard 3D vision transformers scale poorly with image size. Additionally, M3D-LaMed’s vision-language alignment relies on CLIP’s contrastive loss, which compares each image-text pair against a large set of negative pairs in the batch. This approach works well for large datasets but can be less effective in smaller medical datasets, where meaningful negative samples are limited, and similar images may still share semantic information. Furthermore, its multi-modal fusion mechanism, which projects visual features into the LLM via a multilayer perceptron (MLP), may not sufficiently capture complex cross-modal interactions.
To overcome the limitations of existing VLMs in 3D imaging, we introduce Med3DVLM, a novel VLM that incorporates three key innovations to improve 3D feature learning, cross-modal alignment, and multi-modal projection:
  1. 1.
    Efficient 3D Feature Encoding. We integrate DCFormer[1], a 3D image encoder that efficiently captures volumetric features by decomposing 3D convolutions into three parallel 1D convolutions along the depth, height, and width axes. This approach reduces computational complexity and enables richer, more scalable feature representations for 3D image volumes.
  2. 2.
    Improved Vision-Language Alignment. We adopt SigLIP[42], a sigmoid-based language-image pretraining scheme, to improve image-text alignment. Unlike CLIP’s softmax-based contrastive loss, which relies on distinguishing positive pairs from a large batch of negative samples, SigLIP uses a pairwise sigmoid loss that directly optimizes each image-text pair independently. This eliminates the need for global similarity normalization across batches, making training more stable and less sensitive to batch size.
  3. 3.
    Multi-Scale Multi-Modal Projector. We introduce a novel projector based on MLP-Mixer[29] to fuse image and text embeddings effectively. Utilizing a low-high-hybrid design, it blends detailed low-level and abstract high-level features from the image encoder with LLM text embeddings using stacked MLP layers. Inspired by MLP-Mixer’s ability to mix spatial and feature data, this dual-stream approach captures richer cross-modal interactions than simple linear projection, improving the LLM’s decoding accuracy.
2 Related Work
2.1 Medical Vision-Language Models
Medical VLMs utilize multi-modal learning to enhance tasks such as disease diagnosis, report generation, and medical VQA through improved integration and understanding of medical images and text. One of the earliest efforts, ConVIRT [44], applied contrastive learning to learn visual representations from paried image-text data, outperforming ImageNet pretraining in medical image classification and zero-shot retrieval tasks. Similarly, BioViL [4] leveraged large-scale biomedical datasets to refine multi-modal representations, significantly outperforming previous supervised methods. Further extending these capabilities, MedCLIP [35] replaced the InfoNCE loss with semantic matching loss based on medical knowledge, demonstrating its ability to learn generalized representations with limited data.
Recent efforts have aimed to develop VLMs for 3D medical imaging. RadFM [38] introduced a generalist VLM trained on massive datasets that consisted of both 2D and 3D medical images and associated radiology reports. It integrated contrastive learning and generative modeling across diverse imaging modalities (X-ray, CT, MRI) and textual reports, enabling a unified radiology representation. Extending this, M3D [2] introduced a large-scale dataset, M3D-Data, along with M3D-LaMed, a multi-modal vision-language model for 3D medical image analysis tasks such as image-text retrieval, report generation, VQA, and promotable segmentation. Its successor, E3D-GPT [14], employed a 3D multi-modal masked autoencoder framework to further enhance image-text retrieval, leading to improved performance in report generation and VQA.
2.2 Radiology Report Generation
Automated report generation aims to produce descriptive, accurate, and clinically relevant reports from medical images. It improves diagnostic accuracy while reducing radiologists’ workload and supporting care in resource-limited settings. Early approaches primarily adopted an encoder-decoder architecture, where convolutional neural networks (CNNs) extracted image features, and long short-term memory (LSTM) networks generated text descriptions [39, 12, 41, 36]. However, these methods struggled with long-range dependencies, often produced repetitive text, and had limited capacity to capture complex medical semantics.
The transformer architecture [32] addressed key limitations of CNNs and LSTMs by using self-attention to capture global dependencies. Pretrained LLMs [30, 26] based on transformers are now adapted into medical VLMs for report generation. R2GenGPT [34] introduced an efficient visual alignment module to better integrate image features with LLM word embeddings, improving text coherence and clinical relevance. Med-Flamingo [22], based on OpenFlamingo-9B, was pretrained on interleaved medical image-report pairs, achieving superior performance in generating clinically useful responses. More recent methods, such as LLaVA-Med, were based on the LLaVA [19] framework, which combined a visual encoder with an LLM to generate detailed and coherent reports. CT2Rep [11] proposed an advanced 3D vision encoder to generate radiology reports for 3D medical imaging, specifically targeting chest CT volumes. Similarly, CT-CHAT [10] adapted the LLaVA framework for chest CT report generation, demonstrating the effectiveness of large-scale pretrained LLMs in 3D medical imaging.
2.3 Medical Visual Question Answering
A medical VQA system can answer natural language questions about medical images. This task can be closed-ended (with answers like yes/no or a choice from a fixed list) or open-ended (free-form text answers). Early work on medical VQA largely treated it as a classification problem, where models selected the correct answer from a predefined set of possible responses [23, 45, 9, 18]. These models typically encoded images (e.g., using CNNs) and text (e.g., using transformers), and then mapped the combined features to a predefined answer space. Although this approach worked well for simple questions, such as identifying an organ, it struggled with questions requiring detailed explanation or answers not included in the predefined list.
Nowadays, approaches to medical VQA increasingly adopt generative architectures, treating answer generation as a sequence prediction task, which enables open-ended responses. An early step in this direction was CGMVQA [28], which added a generative decoder branch alongside the traditional classifier. Recent VQA systems have fully transitioned to generative models [7]. These models (e.g., Med-PaLM M [31] and LLaVA-Med [15]) extend an LLM with visual inputs via visual instruction tuning [19], allowing the LLM to interpret medical images and answer questions in a conversational manner. Recently, a few works have explored VQA on 3D medical images [2, 10], where models can directly perceive entire 3D image volumes and generate context-aware responses, offering more accurate and holistic medical insights.
3 Methods
3.1 Dataset
This study used the publicly available M3D dataset [2], collected with informed consent and ethical approval by the original investigators. All data were de-identified, and no additional consent or approval was required for this secondary analysis. We used two subsets of the M3D dataset: M3D-Cap and M3D-VQA. M3D-Cap includes 120K image-text pairs, while M3D-VQA contains 662K instruction-response pairs. The dataset can support tasks such as image-text retrieval, radiology report generation, and VQA. All image volumes were resampled to a fixed size of 128x256×256. The dataset was divided into training (115k samples), validation (3k samples), and test (2k samples) sets.
3.2 Med3DVLM
We introduce Med3DVLM, a 3D medical VLM consisting of three core components: a vision encoder, a multi-modal projector, and a large language model. The vision encoder extracts detailed visual features from 3D medical image volumes. These features are integrated with text embeddings through the multi-modal projector, facilitating cross-modal interactions. The LLM then generates coherent, contextually accurate outputs based on these fused features. The overall model architecture is illustrated in Figure 1. The training process is divided into three stages: (1) contrastive pretraining, (2) multi-modal projector pretraining, and (3) VLM fine-tuning.
Figure 1: Overview of our Med3DVLM model.
3.2.1 Contrastive Pretraining
Following the CLIP framework [27], the vision and text encoders are trained jointly using a contrastive loss to align 3D images with their corresponding radiology reports. The goal is to maximize the similarity between embeddings of positive (matching) image-text pairs while minimizing it for negative (non-matching) pairs.
For the vision encoder, we utilize DCFormer [1], which efficiently extracts volumetric features using decomposed 3D convolutions. This approach mitigates computational challenges in high-resolution 3D image analysis, where traditional 3D CNNs and vision transformers struggle due to the cubic scaling of 3D CNNs with kernel size and the quadratic scaling of vision transformers with image size. DCFormer decomposes 3D convolutions into three 1D parallel depthwise convolutions:
$X^{*}_{\text{h}}$ $=\text{DWConv}^{k_{h}\times 1\times 1}_{C\rightarrow C}(X),$ (1)
$X^{*}_{\text{w}}$ $=\text{DWConv}^{1\times k_{w}\times 1}_{C\rightarrow C}(X),$ (2)
$X^{*}_{\text{d}}$ $=\text{DWConv}^{1\times 1\times k_{d}}_{C\rightarrow C}(X),$ (3)
where $X\in\mathbb{R}^{B\times C\times H\times W\times D}$ denote the input feature map, $B$ is the batch size, $C$ is the number of channels, and $H$, $W$, and $D$ represent the spatial dimensions (height, width, and depth), respectively. The kernel sizes in these dimensions, denoted as $(k_{h},k_{w},k_{d})$, are set by default to $k_{h}=k_{w}=k_{d}=k\in\{13,11,9,7\}$ to leverage large kernels for improved feature extraction. The extracted features are then normalized and summed to produce the final output:
$X^{\prime}$ $=X+\text{Norm}_{h}(X^{*}_{\text{h}})+\text{Norm}_{w}(X^{*}_{\text{w}})+\text{Norm}_{d}(X^{*}_{\text{d}}).$ (4)
Such a decomposition strategy enables the processing of 3D volumes at a larger size of 128x256×256 in this study, preserving fine-grained spatial details for improved image-text alignment. In this work, we utilize DCFormer-small variant, which consists of a stem stage and four hiararchical stages with [2, 3, 6, 2] layers and [96, 192, 384, 768] channels [1].
For the text encoder, we use ClinicalBERT [33], an adaptation of BERT [6] pretrained on clinical notes. Designed for clinical text, ClinicalBERT is well-suited for processing radiology reports, generating embeddings that effectively capture their semantic content.
Traditional softmax-based contrastive learning (e.g., CLIP) distinguishes positive pairs from a large batch of negative samples. While effective for large datasets, this approach struggles with smaller medical datasets, where meaningful negative samples are limited, and similar images may share semantic information. To address this, we adopt SigLIP [42], a sigmoid-based language-image pretraining scheme that directly optimizes each image-text pair independently. This eliminates the need for global similarity normalization across batches, resulting in more stable training and reduced sensitivity to batch size. Sigmoid loss is defined as:
$\mathcal{L}_{\text{SigLIP}}=-\frac{1}{|B|}\sum_{i=1}^{|B|}\sum_{j=1}^{|\mathcal{B}|}\log\underbrace{\frac{1}{1+e^{z_{ij}(-t\mathbf{x}_{i}\cdot\mathbf{y}_{j}+b)}}}_{\mathcal{L}_{ij}},$ (5)
where $z_{ij}$ is the label for the $i$-th image and $j$-th text pair, $z_{ij}=1$ for positive pairs and $z_{ij}=0$ otherwise. The sigmoid loss $\mathcal{L}_{ij}$ is computed for each image-text pair, with $t$ as the temperature parameter, $\mathbf{x}_{i}$ and $\mathbf{y}_{j}$ as the image and text embeddings, respectively, and $b$ as a bias term. The overall loss is averaged over all pairs in the batch.
3.2.2 Multi-modal Projector Pretraining
We propose an effective multi-modal projector for the interaction between visual and textual embeddings. In this pretraining stage, all model weights are frozen except for those in the projector. To effectively capture spatial and feature dependencies, we adopt the MLP-Mixer architecture [29], which alternates between a token-mixing MLP (spatial mixer) and a channel-mixing MLP. To the best of our knowledge, this is the first use of the MLP-Mixer in 3D VLMs.
Given an input image feature map $X\in\mathbb{R}^{n\times d}$, where $n$ is the number of tokens and $d$ is the embedding dimension, the MLP-Mixer operations are defined as:
$U$ $=W_{2}\,\sigma(W_{1}\,(\text{Norm}(X)^{\top})),$ (6)
$Y$ $=W_{4}\,\sigma(W_{3}\,\text{Norm}(U^{\top})),$ (7)
where $W_{1},W_{2},W_{3},W_{4}$ are learnable weight matrices, and $\sigma$ denotes a nonlinear activation function (e.g., GELU). The token-mixing MLP captures interactions across spatial tokens, while the channel-mixing MLP models dependencies across feature channels.
To further enhance multi-modal fusion, we introduce a low-high hybrid Mixer-MLP architecture (Figure 2), inspired by Janus Pro [5]. Unlike M3D-LaMed [2], which produces large spatial outputs $(B,2048,768)$ that require lossy downsampling, DCFormer outputs semantically rich features $(B,32,768)$ from its final layer and spatially detailed features $(B,256,384)$ from its penultimate layer. To leverage both types of features, we utilize two parallel MLP-Mixer modules—one for each layer. After passing through $N$ Mixer layers, the resulting feature outputs are concatenated into a unified sequence of image tokens, which are then combined with text tokens from the tokenizer and fed into the LLM.
This hybrid design enables the projector to fuse high-level abstract semantics with low-level spatial details, improving multi-modal alignment. Unlike prior approaches that downsample features at the expense of context, our method preserves diverse information without increasing token length. As a result, our projector produces richer and more semantically aligned joint representations, significantly enhancing radiology report generation and VQA.
Figure 2: Multi-modal projector based on multi-scale MLP-Mixer.
3.2.3 VLM Fine-tuning
In this stage, we fine-tune the multi-modal projector and the LLM for the report generation and VQA. Instead of fine-tuning the entire LLM, we adopt Low-Rank Adaptation (LoRA) [13], which significantly reduces the number of trainable parameters by introducing low-rank updates to the model’s weights. This approach enables efficient adaptation while preserving the LLM’s general knowledge.
Specifically, LoRA modifies a weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ by adding a learnable low-rank update:
$W=W_{0}+\Delta W,\quad\text{where}\quad\Delta W=BA$ (8)
where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ are low-rank matrices with $r\ll\min(d,k)$. This decomposition allows task-specific tuning using far fewer parameters than full fine-tuning, resulting in faster training and lower memory usage.
During fine-tuning, only the weights of the multi-modal projector and the LoRA modules are updated, while all other parameters of the LLM remain frozen. The LLM is initialized with Qwen2.5-7B-Instruct [40], a model pretrained for instruction-following tasks. By jointly optimizing the LoRA parameters and the multi-modal projector, our model effectively adapts to the characteristics of the M3D-Cap and M3D-VQA datasets.
3.3 Evaluation Metrics
We evaluate image-text retrieval performance using Recall@K (R@K) metrics, which quantify the percentage of correct matches within the top K retrieved items. Specifically, R@1 represents the proportion of queries where the correct match is ranked first, while R@5 and R@10 indicate the proportion of queries where the correct match appears within the top 5 or 10 results, respectively.
We evaluate the quality of radiology report generation and open-ended VQA using four key metrics: BLEU [24], ROUGE [16], METEOR [3], and BERTScore [43]. BLEU measures precision by evaluating the overlap of n-grams between the generated text and the reference text, focusing on how many words or phrases in the generated report are found in the reference text. ROUGE emphasizes recall, assessing how much of the reference text appears in the generated output, which is particularly useful for summarization tasks to ensure key details are retained. METEOR improves upon BLEU and ROUGE by considering synonyms and stemming, balancing precision and recall to provide a more flexible and comprehensive measure. BERTScore utilizes deep learning-based contextual embeddings to evaluate semantic similarity.
3.4 Implementation Details
We implemented Med3DVLM using PyTorch [25] and Hugging Face Transformers [37], leveraging DeepSpeed ZeRO2 and BF16 precision on 8 NVIDIA A100 80GB GPUs. We used the AdamW optimizer [20] with gradient accumulation steps of 1, a warmup ratio of 0.03, a cosine learning rate scheduler, and weight decay where specified.
During contrastive pretraining on the M3D-Cap dataset, we resized 3D CT volumes to $128\times 256\times 256$, used DCFormer-small [1] as the vision encoder, and ClinicalBERT [33] as the text encoder. Image features were aggregated via mean pooling, while text features were derived from the [CLS] token, as it captures the global semantic representation of the entire sequence in ClinicalBERT, making it suitable for contrastive alignment with image features. Both were projected to a shared 768-dimensional space. We used a batch size of 64, learning rate of $1\times 10^{-4}$, weight decay of 0.1, and ran for 100 epochs.
For VLM pretraining, we introduced a hybrid MLP-Mixer projector, which fused low-level features from the penultimate encoder layer ($256\times 384$) and high-level features from the final layer ($32\times 768$). Each feature stream passed through $N$ Mixer layers, and the resulting tokens were concatenated and paired with text tokens as input to the language model. We excluded yes/no questions from M3D-VQA to reduce binary-response bias. This stage used a batch size of 16, learning rate of $1\times 10^{-4}$, no weight decay, and was trained for 3 epochs.
Finally, during VLM fine-tuning, we used LoRA [13] with a rank of 16, scaling factor $\alpha=32$, and dropout of 0.05, updating only the LoRA modules and the projector while keeping all other weights frozen. We fine-tuned the model on both M3D-Cap and M3D-VQA (including yes/no questions), using a batch size of 8, learning rate of $5\times 10^{-5}$, no weight decay, and training for 5 epochs. All M3D-LaMed results are reported as in the original paper [2].
4 Results
4.1 Image-Text Retrieval
Med3DVLM significantly outperforms M3D-LaMed in image-to-text and text-to-image retrieval tasks, as shown in Table 1. It achieves the highest Recall@1 across all test set sizes (100, 500, 1000, and 2000). These consistent gains highlight its ability to learn semantically rich representations from high-resolution volumetric data and align them with corresponding text descriptions.
Notably, the performance gap between Med3DVLM and M3D-LaMed widens with larger test sets, suggesting that Med3DVLM generalizes better under more challenging and diverse retrieval scenarios. This trend indicates robustness not only to variation in anatomical content and imaging conditions but also to increasing data scale—a critical requirement for deployment in real-world clinical settings. In such environments, retrieval systems must support large-scale databases while maintaining high precision.
Table 1: Image-text retrieval performance on the M3D dataset. R@K: Recall@Top K. IR: image retrieval. TR: text retrieval. DCFormer-S: small version of DCFormer.
Methods PMC-CLIP M3D-LaMed DCFormer-S SigLIP (Ours)
Test Samples 100 500 1000 2000 100 500 1000 2000 100 500 1000 2000
IR R@1 9.00 4.40 1.90 1.15 64.00 39.60 27.30 19.10 92.00 77.20 69.30 61.00
R@5 28.00 12.80 7.60 4.35 95.00 76.20 61.10 47.45 95.00 92.80 91.40 87.20
R@10 45.00 18.80 12.10 7.60 99.00 87.20 76.10 62.25 96.00 95.20 93.90 91.25
TR R@1 18.00 7.60 4.60 3.15 70.00 40.40 26.60 18.45 90.00 78.40 71.30 63.90
R@5 47.00 20.20 13.00 8.55 95.00 74.20 61.80 47.30 95.00 93.40 91.40 88.05
R@10 59.00 31.00 19.80 13.55 98.00 87.00 75.30 62.15 95.00 95.60 94.10 91.90
4.2 Radiology Report Generation
4.2.1 Quantitative Performance Evaluation
Med3DVLM achieves state-of-the-art performance in radiology report generation, as shown in Table 2. It outperforms all baselines across BLEU, ROUGE, and METEOR, with particularly notable gains in METEOR, which captures both fluency and content relevance. These results suggest that Med3DVLM can generate clinically coherent, well-structured descriptions of complex imaging findings.
Interestingly, although Med3DVLM significantly outperforms RadFM and M3D-LaMed on n-gram-based metrics, all three methods achieve comparable BERTScores. This suggests that while the generated texts may be similarly aligned with reference reports at a high semantic level, Med3DVLM produces outputs with more precise lexical choices, better grammar, and clearer structure. The discrepancy between BERTScore and traditional metrics highlights the importance of evaluating both semantic fidelity and linguistic quality—especially in medical reporting, where accuracy and clinical readability are essential.
Table 2: Report generation performance on the M3D-Cap dataset.
Method BLEU ROUGE METEOR BERTScore
RadFM 12.23 16.49 11.57 87.93
M3D-LaMed (Linear) 14.49 19.25 14.11 88.32
M3D-LaMed (MLP) 15.15 19.55 14.38 88.46
Med3DVLM (Ours) 36.88 40.25 36.42 88.11
4.2.2 Qualitative Analysis of Generated Reports
Figure 3 compares the radiology reports generated by Med3DVLM and M3D-LaMed on a chest CT. Med3DVLM correctly identifies key abnormalities, including multifocal hepatic mass lesions, portal vein thrombosis, and occlusive filling defects, aligning well with the ground truth. It also captures scattered calcifications and small-volume pelvic free fluid, findings not explicitly mentioned but plausibly related to the documented pathology. In contrast, M3D-LaMed fails to detect any liver pathology, underscoring its limitations in aligning imaging features with diagnostic text.
Despite its improvements, Med3DVLM generates incorrect findings, such as hepatic steatosis, contrast enhancement, and iliac arteriovenous malformation, suggesting a tendency to overgeneralize common radiological patterns. Mentions of left ureter abnormalities and renal cortical thinning appear to be hallucinated, indicating the need for better factual grounding. M3D-LaMed, while not generating outright hallucinations, produces findings entirely unrelated to liver pathology, such as bilateral ovarian vein enlargement and para-uterine venous varicosities, suggesting weaker cross-modal alignment.
Figure 3: An example of generated radiology reports by Med3DVLM and M3D-LaMed. Med3DVLM demonstrates improved alignment with clinical findings, while M3D-LaMed generates irrelevant or incorrect content.
4.3 Visual Question Answering
4.3.1 Quantitative Performance Evaluation
Med3DVLM achieves state-of-the-art performance in VQA, as shown in Tables 3 and 4. It consistently outperforms all baselines across both open-ended and closed-ended tasks, demonstrating superior capability in interpreting and reasoning over 3D medical images. The model achieves particularly strong results in categories such as organ and abnormality identification, reflecting its ability to extract detailed anatomical and contextual information.
Med3DVLM demonstrates greater improvements in open-ended VQA than in closed-ended tasks (9.5% in METEOR vs. 5.2% in accuracy), highlighting its strength in generating clinically relevant free-form responses that require a deeper understanding of medical imaging and context. Its strong METEOR performance suggests that its responses are not only well-aligned with ground truth but also more coherent and fluently structured than those from RadFM and M3D-LaMed. In contrast, the smaller performance gap in closed-ended VQA suggests that classification-based answering relies more on direct pattern recognition than on the model’s ability to generate expressive and context-aware responses.
Table 3: Open-ended visual question answering performance on the M3D-VQA dataset.
Open-ended VQA
Method Metric Plane Phase Organ Abnormality Location Mean
RadFM BLEU 14.24 14.25 14.24 15.64 23.58 16.39
ROUGE 25.40 25.41 25.38 25.38 29.09 26.13
METEOR 20.62 20.63 20.61 20.60 24.19 21.33
BERTScore 92.68 92.04 86.79 85.84 86.26 88.72
M3D-LaMed BLEU 98.37 74.41 34.20 15.91 24.00 49.38
ROUGE 98.42 78.63 37.87 19.27 27.74 52.39
METEOR 49.20 63.58 23.78 12.83 18.50 33.58
BERTScore 99.47 95.55 88.97 86.08 87.60 91.53
Med3DVLM (Ours) BLEU 98.85 78.17 40.22 18.99 25.66 52.38
ROUGE 98.89 84.20 45.22 23.27 29.99 56.31
METEOR 49.43 68.50 29.32 16.21 20.32 36.76
BERTScore 99.83 96.47 90.47 86.27 87.88 92.18
Table 4: Comparison of closed-ended visual question answering performance with state-of-the-art methods on the M3D-VQA dataset. The results are reported in terms of accuracy.
Close-ended VQA
Methods Plane Phase Organ Abnormality Location Mean
RadFM 19.65 28.70 16.80 18.92 14.88 19.79
M3D-LaMed 98.80 79.75 74.75 66.65 58.94 75.78
Med3DVLM (Ours) 99.15 87.50 77.45 70.17 64.49 79.75
4.3.2 Qualitative Analysis of VQA Accuracy
Figure 4 presents a qualitative comparison of Med3DVLM and M3D-LaMed in open-ended and closed-ended VQA. In open-ended VQA, both M3D-LaMed and Med3DVLM correctly recognizes the axial imaging plane and localizes a mass lesion in the liver. However, in more complex cases, such as lesion localization in the cranial fossa, Med3DVLM provides a partially correct answer, whereas M3D-LaMed misidentifies the region entirely. While Med3DVLM significantly improves answer accuracy, errors in lesion localization indicate room for enhancement in fine-grained spatial reasoning and uncertainty calibration.
In closed-ended VQA, Med3DVLM consistently outperforms M3D-LaMed by correctly identifying the pleural effusion in the right lung, the mass lesion in the stomach, and the portal venous phase of the CT scan. In contrast, M3D-LaMed misclassifies the pleural effusion as being in the left lung, the stomach lesion as adenocarcinoma, and the CT phase as contrast phase. These errors indicate M3D-LaMed’s limitations in recognizing imaging features and aligning them with clinical knowledge.
Figure 4: Examples of open-ended and closed-ended VQA results by Med3DVLM and M3D-LaMed.
4.4 Ablation Study
4.4.1 Impact of Vision Encoder and Contrastive Loss
We evaluated the effect of different vision encoders (ViT3D vs. DCFormer) and contrastive loss functions (CLIP vs. SigLIP) on image-text retrieval performance. The results in Table 5 indicate that DCFormer significantly outperforms ViT3D, particularly at larger test set sizes, confirming its effectiveness in preserving spatial and structural information in 3D medical images. Additionally, switching from CLIP to SigLIP further improves retrieval performance, where Med3DVLM achieves the highest recall across all test sizes. These findings suggest that SigLIP better handles small-batch medical datasets by improving semantic alignment and feature stability in contrastive learning.
Table 5: The impact of vision encodes and loss functions on image-text retrieval. DCFormer-S: small version of DCFormer.
Methods ViT3D CLIP DCFormer CLIP DCFormer-S SigLIP (Ours)
Test Samples 100 500 1000 2000 100 500 1000 2000 100 500 1000 2000
IR R@1 63.00 33.40 25.90 19.10 84.00 52.00 41.10 31.20 92.00 77.20 69.30 61.00
R@5 89.00 67.80 57.20 44.70 95.00 85.40 77.20 67.15 95.00 92.80 91.40 87.20
R@10 95.00 81.00 71.60 59.35 97.00 90.80 85.90 78.35 96.00 95.20 93.90 91.25
TR R@1 57.00 32.00 25.50 19.40 81.00 53.20 41.10 30.35 90.00 78.40 71.30 63.90
R@5 90.00 67.40 56.90 45.35 95.00 84.80 76.40 64.70 95.00 93.40 91.40 88.05
R@10 94.00 81.80 70.70 59.10 98.00 91.80 86.00 78.20 95.00 95.60 94.10 91.90
Table 6: The impact of multi-modal projectors on radiology report generation. All models use DCFormer-small and the Qwen-2.5-7B-Instruct LLM. H: low high hybrid.
Method BLEU ROUGE METEOR BERTScore
2xMLP 15.63 19.86 15.10 84.24
2xMLP-H 15.99 20.25 15.68 84.31
1xMLP-Mixer-H 23.93 27.70 23.25 85.73
2xMLP-Mixer-H 36.88 40.25 36.42 88.11
4.5 Impact of Multi-Modal Projectors
We compared the performance of our model configured with different multi-modal projectors. The results in Table 6 show that DCFormer with 2×MLP-Mixer achieves the highest BLEU, ROUGE, and METEOR scores, confirming that this architecture enables better representation fusion for text generation. In contrast, replacing the 2×MLP-Mixer with simpler MLP-based projection layers leads to a significant drop in fluency and content relevance, as evidenced by lower METEOR scores. These results highlight the importance of low-high feature integration for structured radiology report generation.
Additionally, we investigated the role of multi-modal projectors in both open-ended and closed-ended VQA. In open-ended VQA, where generating clinically relevant responses requires deeper contextual understanding, the 2×MLP-Mixer projector achieves the highest BLEU, ROUGE, and METEOR scores (Table 7). Meanwhile, models using solely MLP-based projectors generate reasonable responses but show limitations in coherence and fluency, reinforcing the importance of multi-level feature aggregation in medical VQA. For closed-ended VQA, the performance gap between projectors is smaller, suggesting that classification-based answering relies more on direct pattern recognition and dataset priors rather than feature expressiveness. Nonetheless, the 2×MLP-Mixer maintains the highest overall accuracy, indicating that richer feature fusion still contributes to more precise and consistent decision-making.
Table 7: The impact of multi-modal projectors on open-ended visual question answering. All models used the DCFormer-small vision encoder and Qwen 2.5-7B-Instruct model. H: low high hybrid.
Open-ended VQA
Method Metric Plane Phase Organ Abnormality Location Mean
2xMLP BLEU 98.67 74.23 33.95 16.92 23.79 49.51
ROUGE 98.72 80.72 38.46 20.97 28.04 53.38
METEOR 49.35 64.69 24.30 14.63 18.97 34.39
BERTScore 99.80 95.75 89.50 85.88 87.64 91.71
2xMLP-H BLEU 98.86 74.46 34.37 16.96 24.49 49.83
ROUGE 98.91 80.87 38.55 20.82 28.81 53.59
METEOR 49.48 64.88 24.42 14.36 19.29 34.49
BERTScore 99.83 95.75 89.50 85.88 87.80 91.75
1xMLP-Mixer-H BLEU 98.85 76.37 38.87 17.47 24.62 51.24
ROUGE 98.93 82.69 42.93 21.46 28.94 54.99
METEOR 49.51 67.00 26.86 15.19 19.63 35.64
BERTScore 99.83 96.16 90.29 85.99 87.80 92.01
2xMLP-Mixer-H BLEU 98.85 78.17 40.22 18.99 25.66 52.38
ROUGE 98.89 84.20 45.22 23.27 29.99 56.31
METEOR 49.43 68.50 29.32 16.21 20.32 36.76
BERTScore 99.83 96.47 90.47 86.27 87.88 92.18
Table 8: The impact of multi-modal projectors on closed-ended visual question answering. All models used the DCFormer-small vision encoder and Qwen 2.5-7B-Instruct LLM. H: low high hybrid.
Close-ended VQA
Methods Plane Phase Organ Abnormality Location Mean
2xMLP 98.75 84.35 75.90 69.55 64.70 78.65
2xMLP-H 98.75 86.40 75.05 69.90 62.67 78.55
1xMLP-Mixer-H 99.30 86.10 77.05 70.19 63.92 79.31
2xMLP-Mixer-H 99.15 87.50 77.45 70.17 64.49 79.75
5 Discussion
5.1 Addressing Limitations of Existing Models
While previous 3D VLMs such as M3D-LaMed and RadFM have made significant advances, they remain constrained by computational inefficiency, limited capacity for fine-grained spatial reasoning, and suboptimal multi-modal alignment. Med3DVLM addresses these gaps through three targeted innovations. First, Med3DVLM uses DCFormer to decompose 3D convolutions to reduce complexity while preserving spatial detail, allowing for the processing of high-resolution image volumes. Second, SigLIP improves contrastive learning on small, semantically dense medical datasets by avoiding reliance on large negative batches. Third, the dual-stream MLP-Mixer projector effectively fuses low- and high-level image features with text embeddings, enriching the semantic alignment with the LLM.
5.2 Clinical Implications
The integration of Med3DVLM into clinical workflows offers strong potential to improve medical imaging interpretation, decision support, and patient care. Its robust performance in image-text retrieval enables accurate alignment between 3D imaging studies and corresponding textual reports, facilitating rapid access to similar prior cases. This capability supports comparative analysis, longitudinal tracking, and second-opinion workflows, and can enhance diagnostic accuracy and consistency. In educational settings, it may also serve as a valuable tool for case-based learning and clinical training. Med3DVLM also generates accurate, fluent radiology reports directly from 3D volumetric data, reducing documentation burden—especially in high-volume clinical environments. Its substantial gains in METEOR and ROUGE scores indicate strong alignment with expert-level interpretations, making it a practical tool for assisting in preliminary report drafting and standardizing report quality. Finally, the model’s VQA functionality allows clinicians to interact with 3D scans using natural language. This enables task-specific, context-aware responses that support anatomical identification, diagnostic clarification, and real-time clinical decision-making—ultimately reducing the risk of oversight and enhancing confidence in image interpretation.
5.3 Challenges and Future Directions
While Med3DVLM sets a new standard in 3D VLM, several challenges remain. The model occasionally generates hallucinated content in radiology reports, such as references to unrelated anatomical findings, highlighting the persistent issue of factual grounding in generative VLMs. Additionally, although performance in open-ended VQA improved markedly, some errors in lesion localization and answer precision persist, suggesting the need for enhanced spatial reasoning and more robust alignment between visual features and textual outputs.
As future directions, incorporating structured clinical knowledge could help mitigate hallucinations and improve interpretability. Validating Med3DVLM across diverse clinical datasets and imaging protocols will be essential to ensure its reliability in real-world settings. Moreover, extending the model to support uncertainty-aware response generation may further enhance its trustworthiness in clinical decision support.
6 Conclusion
We presented Med3DVLM, a vision-language model designed specifically for 3D medical image analysis. Med3DVLM integrates an efficient volumetric encoder (DCFormer), a sigmoid-based contrastive learning strategy (SigLIP), and a dual-stream MLP-Mixer projector. Extensive experiments on the M3D dataset demonstrate that Med3DVLM achieves state-of-the-art performance in image-text retrieval, radiology report generation, and visual question answering. Notably, Med3DVLM maintains high accuracy while remaining computationally efficient, making it suitable for deployment in real-world clinical workflows. These results underscore the promise of 3D VLMs as foundational tools for building automated, interpretable, and multimodal medical AI systems.